A 50 loop circular coil has a radius of 3cm .it is oriented so that the field lines of a magnetic field are normal to the area of the coil .if the magnetic field is varied so that vector B increase from 0.10T to 0.35T in 2milliseconds,what is the average induced emf in the coil?
1
Expert's answer
2020-05-20T09:08:59-0400
"\epsilon=NA\frac{B_f-B_i}{t}=\pi r^2N\frac{B_f-B_i}{t}"

"\epsilon=\pi (0.03)^2(50)\frac{0.35-0.1}{0.002}=18\ V"